OpenStudy (anonymous):

simplify? -8(3y-2)-(2y+6)-12(y+1) is it 10y-11?

OpenStudy (anonymous):

i multiplied -8 to 3y & 2 which is -8*3y=24 -8*2=16

OpenStudy (anonymous):

then 12 * y 12* 1

OpenStudy (lilg132):

oops made a mistake

OpenStudy (lilg132):

yes one second

OpenStudy (anonymous):

ok

OpenStudy (lilg132):

-24y + 16 -2y -6 -12y -12 -24y - 2y - 12y - 12 + 16 - 6 -38y -2

OpenStudy (lilg132):

tell me if you dont understand something

OpenStudy (anonymous):

how did you get -38?

OpenStudy (anonymous):

its not -34?

OpenStudy (anonymous):

oh wait ur suppose to add?

OpenStudy (lilg132):

-8 * 3y = -24y - * 2y = -2y -12 * y = -12y -24y - 2y - 12y = -38y

OpenStudy (lilg132):

if you do on a calculater -24 -2 it will give you - 26

OpenStudy (anonymous):

how do i know when im suppose to add or subtract? that gets me really confused cuz it has a subracting sign but then we add?

OpenStudy (lilg132):

its pretty much the same as adding as long as you remember to put the minus sign before it

OpenStudy (lilg132):

ok well let me try and show you something that might help you

OpenStudy (lilg132):

when you look at that if your on lets say -2 on the number line and you had to do -4 which is the same as -2 -4 you move 4 places to the left which gives you -6 that should help you understand negative numbers better

OpenStudy (anonymous):

yess!

OpenStudy (lilg132):

now if we were on -2 on the number line and we had to do +4 which is the same as -2 + 4 or 4 - 2 and you were on -2 on the number line you would move 4 places to the right which would give you +2 or 2

OpenStudy (anonymous):

i have another question can YOU help me with it? but ima try to answer it & can you let me know if im doing it correctly?

OpenStudy (lilg132):

sure open another question

OpenStudy (anonymous):

here? or another box?